OpenStudy (anonymous):

3x+y=5 and 2x-5y=-2 using elimination

OpenStudy (aripotta):

multiply the first equation by 5

OpenStudy (aripotta):

the first equation is 3x + y = 5, right?

OpenStudy (aripotta):

you have a typo saying it's 3z

OpenStudy (aripotta):

yea, multiply the first equation by 5, what do you get?

OpenStudy (anonymous):

can u show me what ur talking about 3x+y=5 2x-5y=-2

OpenStudy (aripotta):